# Modeling Developer Burnout with GenAI Adoption

### v1

@@ 6. Influence of Developer Characteristics (RQ2)
To answer RQ2, we continue to follow the mixed-methods concurrent embedded approach (Greene, 2007; Morse, 2016). We first used ordinary least squares (OLS) regression models (Fox, 2015) to assess the relationship between developer characteristics and job demands, resources, and burnout. We then qualitatively analyzed open-ended responses on job resources to contextualize our analysis results, surfacing how learning resources are (or are not) made available in practice.
@@ 6.1. Data Analysis
Dependent Variables. To create a composite score of the dependent variables that were measured using multiple indicators (e.g., burnout includes four indicators), we use the outer weights of each indicator estimated in the measurement model (see Table 6 in (Anonym, 2025)). For example, $\text{Burnout}=\sum_{i=1}^{4}\text{Burn}_{i}\cdot w_{i}$, where $\text{Burn}_{i}$ are the items (Burn1–Burn4) and $w_{i}$ are their corresponding outer weights. This approach is standard in survey-based research, where responses to multiple Likert-scale items are aggregated to form a single interval-level measure (Sullivan and Artino Jr, 2013; Chikezie et al., 2025).
@@
Model Assessments. We assessed the models’ explanatory power by calculating the $R^{2}$ and adjusted $R^{2}$ values. In social science research, $R^{2}$ values in the range of 0.10–0.30 are often considered acceptable given the complexity of human behavior (Ozili, 2023). Our results fall within this range.
To address Type I error inflation from multiple comparisons, we applied Benjamini-Hochberg adjustment to the $p$-values (Benjamini and Hochberg, 1995). Table 4 shows the regression analysis results with adjusted $p$-values.
6.2. Results
Table 4. Estimated regression coefficients ($\beta$) for the effects of developer characteristics (Coder, Organization Size (Size), Industry Seniority (Years)) on dependent constructs (Burnout, Organizational Pressure, Workload, Autonomy, Learning Resources)
@@ 6.2. Results
Job Demands. Organizational Pressure is positively associated with coding role ($\beta=0.51,p<0.05$) and organization size ($\beta=0.41,p<0.001$) and are statistically significant. These results are consistent with prior work showing that larger organizations impose adoption mandates, thereby heightening the pressure for developers (Bakker and Demerouti, 2007; Scholze and Hecker, 2023). AI-based coding support has seen significant advances and is touted for improving developer productivity (Peng et al., 2023; Paradis et al., 2025; Dohmke et al., 2023); thus, developers in coding-intensive roles likely encounter greater strain with AI adoption (Harness, 2025; Ravn et al., 2022). Industry seniority is not statistically significant. Thus (H4b) is supported only for role and organization size.
@@
Job Resources. We have empirical support for (H4d) organization size and industry seniority. Organization size is negatively associated with autonomy ($\beta=-0.13,p<0.001$), indicating that developers in larger organizations perceive less discretion and decision-making power. This is consistent with research showing that bureaucratic structures and standardized processes often constrain individual influence (Tummers and Bakker, 2021). Industry seniority is positively associated with autonomy ($\beta=0.18,p<0.001$), suggesting that senior developers have greater say over tasks and responsibilities, echoing findings that experience confers both expertise and organizational credibility (Clausen et al., 2019; Hackman and Oldham, 1976). Prior research has shown that large organizations often introduce new tools such as AI through top-down mandates that restrict individual discretion (Scholze and Hecker, 2023). However, senior developers are better positioned to shape adoption practices by leveraging their expertise and credibility (Clausen et al., 2019).
@@
Two authors first conducted open coding to identify preliminary categories of resources. They then iteratively refined these codes to resolve discrepancies and collaboratively construct a final codebook through negotiated agreement (Garrison et al., 2006; Forman and Damschroder, 2007). Using the codebook, two authors independently coded 20% of the responses and compared results to calculate IRR using the Jaccard index(Landis and Koch, 1977), achieving 92% agreement.
@@ 7. Limitations
In empirical studies of heterogeneous domains, theory is crucial for guiding analysis and preventing purely descriptive interpretations (Stol and Fitzgerald, 2018). We adopted the JD–R model as our theoretical lens because it enables systematic analysis of the interplay between job demands, job resources, and practitioner well-being, and it has been widely applied in organizational and SE contexts (Bakker and Demerouti, 2017; Trinkenreich et al., 2024b; Kim et al., 2024). We acknowledge that some aspects of AI adoption may fall outside the scope of JD–R; however, we included an open-ended survey question to ask participants to share additional thoughts about AI adoption. Our qualitative analysis of these responses revealed no additional themes beyond those covered by our survey items.
Our hypotheses in this study propose associations between different constructs in the JD-R model rather than causal relationships, as the present study is a cross-sectional sample study (Stol and Fitzgerald, 2018). For example, participants who were already experiencing higher burnout (e.g., due to upcoming deadlines) might adopt AI differently compared to those under lighter workloads, which may introduce response bias. Other confounding factors, such as trust in AI systems or perceived productivity, may have influenced results but were not directly measured. Therefore, our results should be interpreted as a theoretical starting point, guiding future studies to explore these contextual influences.
When answering RQ1, we employed PLS–SEM. We acknowledge that there are other methods, such as CB–SEM, but we selected PLS–SEM because our JD–R model includes both formative and reflective constructs, which PLS–SEM is better suited to analyze (Hair et al., 2011). Moreover, PLS–SEM offers stronger explanatory power, making it appropriate for investigating developer well-being in the emerging context of GenAI adoption (Hair et al., 2011).
While no single sample can capture the entire global software workforce, our sample of 442 software practitioners from 56 organizations is comparable in size to, or larger than, many empirical studies in software engineering (Russo, 2024; Trinkenreich et al., 2023). More importantly, our participants represent a wide range of company sizes, job roles, and levels of experience, which supports diversity in perspectives and provides a suitable starting point for understanding the associations presented in our model.
@@ 8. Concluding Remarks
The introduction of GenAI to software engineering has generated extraordinary excitement, with industry narratives promising unprecedented productivity gains. Yet, as with earlier waves of technological enthusiasm, the promise of relief is overlaid with forms of work pressure and strain (Khan et al., 2023).
@@
Shift 2: From apprenticeship to private copy–paste. Pre GenAI, newcomers’ learning progression was coupled with social learning practices such as pair programming, code reviews, mentoring, and public Q&A ecosystems (Feng et al., 2022, 2025a). Today, however, practices like Vibe coding and pair programming with AI have emerged as the new trend (Peng et al., 2023; Ray, 2025). While these tools may accelerate certain tasks, they also narrow social networks, reduce mentoring opportunities, and erode the “resources” that historically supported early-career growth. Participants raised concern that this shift could delay skill acquisition, “I fear that juniors will take much longer to get to an experience level that is comparable to what devs with 10+” [P202].
Shift 3: From productivity gains to hidden collaboration costs. Collaboration friction has long been a fixture in software development (Feng et al., 2025b), GenAI use can increase the workload for colleagues, adding to the friction: “reviewing LLM-generated content such as code and docs wastes time, coworkers are (accidentally but carelessly) sabotaging our work by creating [more] work” [P134]. Another participant said: “AI generates many test cases, including verbose ones, which makes code-reviewing more time-consuming” [P25]. While authors may save time by pasting in AI outputs, reviewers inherit the burden of checking quality, correcting errors, and editing down verbosity; “I spend a far greater share of my time editing their writing and checking the validity of their assertions” [P374].
@@ 8.2. Implications For:
Workload Design. Studies abound that extol AI for its ability to improve productivity (Peng et al., 2023; Paradis et al., 2025). However, adoption of AI can create additional downstream workload, as reported by our participants, which can erase the productivity gains. When paired with misplaced management expectations of ever-increasing output, it causes burnout. In fact, one industry survey found that 77% of AI users reported increased workload because of AI adoption (Nawrat, 2024). This is because, while AI reduces some types of work (e.g., automated code or test generation), it creates other types (e.g., in-depth verification, debugging AI-generated code). AI is shifting the type of work from creation towards curation and oversight, which is cognitively demanding and requires in-depth technical expertise, but is harder to track.
@@
In conclusion, our findings show that current hype cycles and mandate-driven adoption of GenAI can intensify job demands, unevenly distribute resources, and contribute to developer burnout. Yet history reminds us that each technological wave, from assembly lines to the automobile to the internet, has reshaped not only what we produce, but how we work. Some practices will fade, and new ones will emerge as the new norm. As participant eloquently summarized: “I think it has a lot of potential to be used as a carefully curated tool in one’s tool belt, like a chisel or a hammer can be for a woodworker but I think trying to force it’s adoption without proper training to enforce such a mindset is going to lead to a large brain drain in developers which will have negative long term ramifications” [P117]. Sustainable development workflows, therefore, depend not only on adopting AI tools but on building the human and organizational capacity to manage their downstream effects.

### v2

@@ 6. Influence of Developer Characteristics (RQ2)
To answer RQ2, we continue to follow the mixed-methods concurrent embedded approach (greene2007mixed; morse2016mixed). We first used ordinary least squares (OLS) regression models (fox2015applied) to assess the relationship between developer characteristics and job demands, resources, and burnout. We then qualitatively analyzed open-ended responses on job resources to contextualize our analysis results, surfacing how learning resources are (or are not) made available in practice.
@@ 6.1. Data Analysis
Dependent Variables. To create a composite score of the dependent variables that were measured using multiple indicators (e.g., burnout includes four indicators), we use the outer weights of each indicator estimated in the measurement model (see Table 6 in (supply)). For example, $\text{Burnout}=\sum_{i=1}^{4}\text{Burn}_{i}\cdot w_{i}$, where $\text{Burn}_{i}$ are the items (Burn1–Burn4) and $w_{i}$ are their corresponding outer weights. This approach is standard in survey-based research, where responses to multiple Likert-scale items are aggregated to form a single interval-level measure (sullivan2013analyzing; chikezie2025measuring).
@@
Model Assessments. We assessed the models’ explanatory power by calculating the $R^{2}$ and adjusted $R^{2}$ values. In social science research, $R^{2}$ values in the range of 0.10–0.30 are often considered acceptable given the complexity of human behavior (ozili2023acceptable). Our results fall within this range.
To address Type I error inflation from multiple comparisons, we applied Benjamini-Hochberg adjustment to the $p$-values (benjamini1995controlling). Table 4 shows the regression analysis results with adjusted $p$-values.
6.2. Results for RQ2
Table 4. Estimated regression coefficients ($\beta$) for the effects of developer characteristics (Coder, Organization Size (Size), Industry Seniority (Years)) on dependent constructs (Burnout, Organizational Pressure, Workload, Autonomy, Learning Resources; $N=442$)
@@ 6.2. Results for RQ2
Job Demands. Organizational Pressure is positively associated with coding role ($\beta=0.51,p<0.05$) and organization size ($\beta=0.41,p<0.001$) and are statistically significant. These results are consistent with prior work showing that larger organizations impose adoption mandates, thereby heightening the pressure for developers (bakker2007job; scholze2023digital). AI-based coding support has seen significant advances and is touted for improving developer productivity (peng2023impact; paradis2025much); thus, developers in coding-intensive roles likely encounter greater strain with AI adoption (harness2025; ravn2022team). Industry seniority is not statistically significant. Thus, (H4b) is supported only for role and organization size.
@@
Job Resources. We have empirical support for (H4d) organization size and industry seniority. Organization size is negatively associated with autonomy ($\beta=-0.13,p<0.001$), indicating that developers in larger organizations perceive less discretion and decision-making power. This is consistent with research showing that bureaucratic structures and standardized processes often constrain individual influence (tummers2021leadership). Industry seniority is positively associated with autonomy ($\beta=0.18,p<0.001$), suggesting that senior developers have greater say over tasks and responsibilities, echoing findings that experience confers both expertise and organizational credibility (clausen2019danish; hackman1976motivation). Prior research has shown that large organizations often introduce new tools such as AI through top-down mandates that restrict individual discretion (scholze2023digital). However, senior developers are better positioned to shape adoption practices by leveraging their expertise and credibility (clausen2019danish).
@@
Learning resource types reported by developers. Six categories of organizational support were identified from 361 open-ended responses. The table also displays the frequency of each category along with distributions by company size and developer seniority (five ordinal levels each).
@@
Two authors first conducted open coding to identify preliminary categories of resources. They then iteratively refined these codes to resolve discrepancies and collaboratively construct a final codebook through negotiated agreement (garrison2006revisiting; forman2007qualitative). Using the codebook, two authors independently coded 20% of the responses and compared results to calculate IRR using the Jaccard index(landis1977measurement), achieving 92% agreement.
@@ 7. Limitations
In empirical studies of heterogeneous domains, theory is crucial for guiding analysis and preventing purely descriptive interpretations (stol2018abc). We adopted the JD–R model as our theoretical lens because it enables systematic analysis of the interplay between job demands, job resources, and practitioner well-being, and it has been widely applied in organizational and SE contexts (bakker2017job; trinkenreich2024predicting). We acknowledge that some aspects of AI adoption may fall outside the scope of JD–R; however, we included an open-ended survey question to ask participants to share additional thoughts about AI adoption. Our qualitative analysis of these responses revealed no additional themes beyond those covered by our survey items.
Our hypotheses in this study propose associations between different constructs in the JD-R model rather than causal relationships, as the present study is a cross-sectional sample study (stol2018abc). For example, participants who were already experiencing higher burnout (e.g., due to upcoming deadlines) might adopt AI differently compared to those under lighter workloads, which may introduce response bias. Other confounding factors, such as trust in AI systems or perceived productivity, may have influenced results but were not directly measured. Therefore, our results should be interpreted as a theoretical starting point, guiding future studies to explore these contextual influences.
When answering RQ1, we employed PLS–SEM. We acknowledge that there are other methods, such as CB–SEM, but we selected PLS–SEM because our JD–R model includes both formative and reflective constructs, which PLS–SEM is better suited to analyze (hair2011pls). Moreover, PLS–SEM offers stronger explanatory power, making it appropriate for investigating developer well-being in the emerging context of GenAI adoption (hair2011pls).
While no single sample can capture the entire global software workforce, our sample of 442 software practitioners from 56 organizations is comparable in size to, or larger than, many empirical studies in software engineering (russo2024navigating; trinkenreich2023belong). More importantly, our participants represent a wide range of company sizes, job roles, and levels of experience, which supports diversity in perspectives and provides a suitable starting point for understanding the associations presented in our model.
@@ 8. Concluding Remarks
The introduction of GenAI to software engineering has generated extraordinary excitement, with industry narratives promising unprecedented productivity gains. Yet, as with earlier waves of technological enthusiasm, the promise of relief is overlaid with forms of work pressure and strain (khan2023excessive).
@@
Shift 2: From apprenticeship to private copy–paste. Pre GenAI, newcomers’ learning progression was coupled with social learning practices such as pair programming, code reviews, mentoring, and public Q&A ecosystems (feng2022case; feng2025multifaceted). Today, however, practices like Vibe coding and pair programming with AI have emerged as the new trend (peng2023impact; ray2025review). While these tools may accelerate certain tasks, they also narrow social networks, reduce mentoring opportunities, and erode the “resources” that historically supported early-career growth. Participants raised concern that this shift could delay skill acquisition, “I fear that juniors will take much longer to get to an experience level that is comparable to what devs with 10+” [P202].
Shift 3: From productivity gains to hidden collaboration costs. Collaboration friction has long been a fixture in software development (feng2025domains), GenAI use can increase the workload for colleagues, adding to the friction: “reviewing LLM-generated content such as code and docs wastes time, coworkers are (accidentally but carelessly) sabotaging our work by creating [more] work” [P134]. Another participant said: “AI generates many test cases, including verbose ones, which makes code-reviewing more time-consuming” [P25]. While authors may save time by pasting in AI outputs, reviewers inherit the burden of checking quality, correcting errors, and editing down verbosity; “I spend a far greater share of my time editing their writing and checking the validity of their assertions” [P374].
@@ 8.2. Implications For:
Workload Design. Studies abound that extol AI for its ability to improve productivity (peng2023impact; paradis2025much). However, adoption of AI can create additional downstream workload, as reported by our participants, which can erase the productivity gains. When paired with misplaced management expectations of ever-increasing output, it causes burnout. In fact, one industry survey found that 77% of AI users reported increased workload because of AI adoption (unleash2024_ai_workload_upwork). This is because, while AI reduces some types of work (e.g., automated code or test generation), it creates other types (e.g., in-depth verification, debugging AI-generated code). AI is shifting the type of work from creation towards curation and oversight, which is cognitively demanding and requires in-depth technical expertise, but is harder to track.
@@
In conclusion, our findings show that current hype cycles and mandate-driven adoption of GenAI can intensify job demands, unevenly distribute resources, and contribute to developer burnout. Yet, history reminds us that each technological wave, from assembly lines to the automobile to the internet, has reshaped not only what we produce, but how we work. Some practices will fade, and new ones will emerge as the new norm. As participant eloquently summarized: “I think it has a lot of potential to be used as a carefully curated tool in one’s tool belt, like a chisel or a hammer can be for a woodworker but I think trying to force it’s adoption without proper training to enforce such a mindset is going to lead to a large brain drain in developers which will have negative long term ramifications” [P117]. Sustainable development workflows, therefore, depend not only on adopting AI tools but on building the human and organizational capacity to manage their downstream effects.
Acknowledgements.
We thank all survey participants for their time and insights. This work was supported by NSF Grant No. 2303043.